If A completes a particular work in 4 days and B the same work in 12 days. How many days will it take if they work together?
Answer by richwmiller(17219)

1/a+1/b=1/x
1/4+1/12=1/x
12/48+4/48=1/x
16/48=1/x
x=48/16
x=3 hours working together
check
x/a+x/b=1
x/4+x/12=1
3/4+3/12=1
0.75+0.25=1
1=1
ok